# Cambridge CB3 0WA, UK

### v1

@@ Abstract
We show that the general charged, rotating black hole in five-dimensional Einstein-Maxwell theory has a singular extremal limit. Only the known analytic solutions with exactly zero charge or zero angular momenta have smooth extremal horizons. When the two angular momenta are not equal and both are scaled to zero, the solution approaches a new family of static, nonspherical extremal black holes. We also consider general black holes in five-dimensional Einstein-Maxwell-Chern-Simons theory, and show that they also have singular extremal limits except for one special value of the coefficient of the Chern-Simons term (the one fixed by supergravity). Combining this with earlier results showing that extremal black holes have singular horizons in four-dimensional general relativity with small higher derivative corrections, and in anti-de Sitter space with perturbed boundary conditions, one sees that smooth extremal horizons are indeed the exception and not the rule.
@@ 1 Introduction
When the curvature scalars remain finite, the metric and vector potential are $C^{0}$ but not $C^{1}$ at the extremal horizon. Since the horizon becomes infinitely far away in spacelike directions, one can define a near-horizon geometry. When $Q\to 0$, it approaches the Myers-Perry near-horizon geometry. It is natural to expect that when the angular momenta vanish, it will approach the 5D Reissner-Nordström near-horizon geometry. However this is incorrect. We find that the solution approaches the more general static $U(1)\times U(1)$ symmetric near-horizon geometries Kunduri:2009ud when $J_{\phi}$ and $J_{\psi}$ both vanish. These metrics involve a squashed $S^{3}$, and were originally interpreted as the near-horizon geometries of a nonrotating extremal black hole in a background electric field. We will show that the same geometries arise (without a background electric field) for an extremal black hole in the limit $J_{\phi},J_{\psi}\to 0$, with the amount of squashing determined by the ratio $J_{\phi}/J_{\psi}$. This implies that there are static, asymptotically flat extremal black holes that are not spherically symmetric!
@@ 3 Near-horizon geometries
$\omega_{\phi}=\frac{a_{\psi}\left(a_{\phi}+a_{\psi}\right)^{2}+\left(a_{\phi}+2a_{\psi}\right)q}{2\sqrt{a_{\phi}a_{\psi}+q}\left[\left(a_{\phi}+a_{\psi}\right)^{2}+q\right]}\,,\quad\omega_{\psi}=\frac{a_{\phi}\left(a_{\phi}+a_{\psi}\right)^{2}+\left(2a_{\phi}+a_{\psi}\right)q}{2\sqrt{a_{\phi}a_{\psi}+q}\left[\left(a_{\phi}+a_{\psi}\right)^{2}+q\right]}\,,\\ \text{and}\quad Q_{\rm NH}=\frac{\sqrt{3}q^{2}}{2\sqrt{a_{\phi}a_{\psi}+q}\left[\left(a_{\phi}+a_{\psi}\right)^{2}+q\right]}\,,$ (14a)
with $q$, $a_{\phi}$ and $a_{\psi}$ being constants that parametrise charge and rotations along the $\phi$ and $\psi$ axis. Furthermore,
$Q_{2}(x)=4Q_{1}(x)=a_{\phi}a_{\psi}+q+a_{\phi}^{2}x^{2}+a_{\psi}^{2}\left(1-x^{2}\right)$ (14b)
$Q_{3}(x)=\frac{4\left(a_{\phi}a_{\psi}+q\right)\left[\left(a_{\phi}+a_{\psi}\right){}^{2}+q\right]^{2}Q_{1}(x)}{Z(x)}$ (14c)
$Q_{4}(x)=\frac{Z(x)}{16Q_{1}(x)^{2}}$ (14d)
$Q_{5}(x)=\frac{4\left(a_{\phi}+b_{\psi}\right)^{2}\left(a_{\phi}b_{\psi}+q\right)Q_{1}(x)-a_{\phi}b_{\psi}q^{2}}{Z(x)}$ (14e)
$Q_{6}(x)=-\frac{\sqrt{3}q}{Z(x)}\left\{4\left(a_{\phi}+a_{\psi}\right)\left(a_{\phi}a_{\psi}+q\right)Q_{1}(x)-a_{\psi}q\left[a_{\psi}\left(a_{\phi}+a_{\psi}\right)+q\right]\right\}$ (14f)
$Q_{7}(x)=-\frac{\sqrt{3}a_{\psi}q}{4Q_{1}(x)}\,,$ (14g)
@@
$Z(x)=\left[a_{\psi}\left(a_{\phi}+a_{\psi}\right)+q\right]^{3}+\left(a_{\phi}-a_{\psi}\right)\left(a_{\phi}+a_{\psi}\right)^{3}\left(a_{\phi}a_{\psi}+q\right)x^{4}\\ +\left[\left(2a_{\phi}-a_{\psi}\right)a_{\psi}^{2}\left(a_{\phi}+a_{\psi}\right)^{3}+\left(4a_{\phi}-a_{\psi}\right)a_{\psi}\left(a_{\phi}+a_{\psi}\right)^{2}q+\left(2a_{\phi}^{2}+2a_{\phi}a_{\psi}-a_{\psi}^{2}\right)q^{2}\right]x^{2}\,.$ (14h)
@@ 3.1 Perturbative in the angular momenta
$\omega_{\phi}=\pm\frac{c_{1}^{2}}{c_{2}^{2}}\omega_{\psi}\,.$ (25)
@@
The above expression was derived using perturbation theory. However, for $\lambda=0$, it turns out to yield the full nonlinear result (with no $\mathcal{O}(\epsilon^{3})$ correction). At this point, we have no justification for why this is the case, as we only know the geometry perturbatively. However, when comparing with the full numerical results in the next sections, we find that the above expression for $\lambda=0$ agrees with the full nonlinear results to the first twenty decimal places, even when $J_{\phi}$ and $J_{\psi}$ are of the same order as $Q^{3/2}$. This agreement is exact only for $\lambda=0$. Indeed, using the known analytic solution with $\lambda=1$ (see Eqs. (14h)), it is a straightforward exercise to show that the expression for $\frac{S}{|Q|^{3/2}}$ given above is valid only to leading order in $J_{\phi}$ and $J_{\psi}$. The above result assumed $J_{\phi}J_{\psi}\geq 0.$ When $J_{\phi}J_{\psi}<0$, we get instead that
@@ 3.3 Fully nonlinear results
which was computed directly from Eqs. (14h), giving reassurance that our numerical code is correct.
@@ 4.1 Perturbative in the angular momenta
$Q_{3}(x,\rho)=\frac{\sigma_{1}(x)}{A_{2}}\left(1+\epsilon\,\rho^{\gamma}\,q_{3}(x)\right)$
$Q_{4}(x,\rho)=\frac{\sigma_{2}(x)}{A_{1}}\left(1+\epsilon\,\rho^{\gamma}\,q_{4}(x)\right)$
@@ 4.3 Fully nonlinear results
We now briefly discuss the $\lambda\neq 0$ case. For $\lambda=1$, all the scaling dimensions become integers. This is to be expected, since a known solution exists and is analytic in $\rho$ when changed to the relevant gauge (see Eq. (14h) for the corresponding near-horizon geometry). For $\lambda>1$ we always find negative scaling dimensions for at least one mode. These connect to the modes generated by ${\rm G}_{3}$ discussed in section 4.1. For $10<\lambda<$ 1 more interesting dynamics takes place. In particular, the orange and green families displayed in Fig. 2 both have negative or complex scaling dimensions, showing that these are not suitable to connect to an asymptotically flat end.
@@ 6 Discussion
Another surprise concerns the limit of vanishing angular momenta. We found that this limit depends on the ratio $J_{\phi}/J_{\psi}$ as both angular momenta are scaled to zero. One approaches the 5D Reissner-Nordström solution only when $J_{\phi}/J_{\psi}=1$. In general, one approaches a new family of static, extremal black holes which are not spherically symmetric. One can view this as a form of spontaneous breaking of spherical symmetry. The limiting black hole is static and asymptotically flat, so there is nothing preventing it from being spherically symmetric. Since the scaling dimensions computed in Appendix A are all integers, these new solutions probably have smooth horizons [^9] The noninteger scaling dimensions computed in Sec. 4.1 correspond to vector modes which are not sourced in the static, asymptotically flat solution.. It is easy to construct nonspherical static extremal black holes when there is more than one horizon (since there is no force between the black holes) and it is known that 5D black holes are not uniquely characterized by their asymptotic charges, but this is different. It is a new form of nonuniqueness for a single extremal 5D black hole.

### v2

@@ Abstract
We show that the general charged, rotating black hole in five-dimensional Einstein-Maxwell theory has a singular extremal limit. Only the known analytic solutions with exactly zero charge or zero angular momenta have smooth extremal horizons. We also consider general black holes in five-dimensional Einstein-Maxwell-Chern-Simons theory, and show that they also have singular extremal limits except for one special value of the coefficient of the Chern-Simons term (the one fixed by supergravity). Combining this with earlier results showing that extremal black holes have singular horizons in four-dimensional general relativity with small higher derivative corrections, and in anti-de Sitter space with perturbed boundary conditions, one sees that smooth extremal horizons are indeed the exception and not the rule.
@@ 1 Introduction
When the curvature scalars remain finite, the metric and vector potential are $C^{0}$ but not $C^{1}$ at the extremal horizon. Since the horizon becomes infinitely far away in spacelike directions, one can define a near-horizon geometry. When $Q\to 0$, it approaches the Myers-Perry near-horizon geometry. It is natural to expect that when the angular momenta vanish, it will approach the 5D Reissner-Nordström near-horizon geometry. However this is incorrect. We find that the solution approaches the more general static $U(1)\times U(1)$ symmetric near-horizon geometries Kunduri:2009ud when $J_{\phi}$ and $J_{\psi}$ both vanish. These metrics involve a squashed $S^{3}$, and were originally interpreted as the near-horizon geometries of a nonrotating extremal black hole in a background electric field. We will show that the same geometries arise (without a background electric field) for an extremal black hole in the limit $J_{\phi},J_{\psi}\to 0$, with the amount of squashing determined by the ratio $J_{\phi}/J_{\psi}$.
@@ 3 Near-horizon geometries
$\omega_{\phi}=\frac{a_{\psi}\left(a_{\phi}+\hat{\delta}\,a_{\psi}\right)^{2}+\left(a_{\phi}+2\,\hat{\delta}\,a_{\psi}\right)q}{2\sqrt{|a_{\phi}a_{\psi}+q|}\left[\left(a_{\phi}+\hat{\delta}\,a_{\psi}\right)^{2}+\hat{\delta}\,q\right]}\hat{\delta}\,,\quad\omega_{\psi}=\frac{a_{\phi}\left(\hat{\delta}\,a_{\phi}+a_{\psi}\right)^{2}+\left(2\hat{\delta}\,a_{\phi}+a_{\psi}\right)q}{2\sqrt{|a_{\phi}a_{\psi}+q|}\left[\left(a_{\phi}+\hat{\delta}\,a_{\psi}\right)^{2}+\hat{\delta}\,q\right]}\hat{\delta}\,,\\ \text{and}\quad Q_{\rm NH}=\frac{\sqrt{3}q^{2}\hat{\delta}}{2\sqrt{|a_{\phi}a_{\psi}+q|}\left[\left(a_{\phi}+\hat{\delta}\,a_{\psi}\right)^{2}+\hat{\delta}\,q\right]}\,,$ (14a)
with $q$, $a_{\phi}$ and $a_{\psi}$ being constants that parametrise charge and rotations along the $\phi$ and $\psi$ axis and
$\hat{\delta}=\left\{\begin{array}[]{c}\phantom{+}1\quad\text{if}\quad a_{\phi}a_{\psi}+q>0\\ -1\quad\text{if}\quad a_{\phi}a_{\psi}+q<0\end{array}\right.\,.$ (14b)
Furthermore,
$Q_{2}(x)=4Q_{1}(x)=|a_{\phi}a_{\psi}+q|+a_{\phi}^{2}x^{2}+a_{\psi}^{2}\left(1-x^{2}\right)$ (14c)
$Q_{3}(x)=\frac{4|a_{\phi}a_{\psi}+q|\left[\left(a_{\phi}+\hat{\delta}\,a_{\psi}\right)^{2}+\hat{\delta}\,q\right]^{2}Q_{1}(x)}{Z(x)}$ (14d)
$Q_{4}(x)=\frac{Z(x)}{Q_{2}(x)^{2}}$ (14e)
$Q_{5}(x)=\frac{4\left(a_{\phi}+\hat{\delta}\,a_{\psi}\right)^{2}\left(a_{\phi}a_{\psi}+q\right)Q_{1}(x)-a_{\phi}a_{\psi}q^{2}}{Z(x)}$ (14f)
$Q_{6}(x)=-\frac{\sqrt{3}q}{Z(x)}\left\{4\left(a_{\phi}+\hat{\delta}\,a_{\psi}\right)|a_{\phi}a_{\psi}+q|Q_{1}(x)-a_{\psi}q\,\hat{\delta}\,\left[a_{\psi}\left(a_{\phi}+\hat{\delta}\,a_{\psi}\right)+q\right]\right\}$ (14g)
$Q_{7}(x)=-\frac{\sqrt{3}a_{\psi}q}{Q_{2}(x)}\,,$ (14h)
@@
$Z(x)=\left(a_{\psi}^{2}+|a_{\phi}a_{\psi}+q|\right)^{3}+\left(a_{\phi}^{2}-a_{\psi}^{2}\right)\left(a_{\phi}+\hat{\delta}\,a_{\psi}\right)^{2}|a_{\phi}a_{\psi}+q|x^{4}\\ +\Big{[}\left(2a_{\phi}-\hat{\delta}\,a_{\psi}\right)a_{\psi}^{2}\left(a_{\phi}+\hat{\delta}\,a_{\psi}\right)^{3}+\left(4a_{\phi}-\hat{\delta}\,a_{\psi}\right)a_{\psi}\left(a_{\phi}+\hat{\delta}\,a_{\psi}\right)^{2}q\\ +\left(2a_{\phi}^{2}+2a_{\phi}a_{\psi}\hat{\delta}-a_{\psi}^{2}\right)q^{2}\Big{]}x^{2}\,.$ (14i)
@@ 3.1 Perturbative in the angular momenta
$\omega_{\phi}=\pm\frac{c_{1}^{2}}{c_{2}^{2}}\omega_{\psi}\,\qquad{\rm for\ all}\ \lambda^{2}\neq 1\,.$ (25)
@@
The above expression was derived using perturbation theory. However, for $\lambda=0$, it turns out to yield the full nonlinear result (with no $\mathcal{O}(\epsilon^{3})$ correction). At this point, we have no justification for why this is the case, as we only know the geometry perturbatively. However, when comparing with the full numerical results in the next sections, we find that the above expression for $\lambda=0$ agrees with the full nonlinear results to the first twenty decimal places, even when $J_{\phi}$ and $J_{\psi}$ are of the same order as $Q^{3/2}$. This agreement is exact only for $\lambda=0$. Indeed, using the known analytic solution with $\lambda=1$ (see Eqs. (14i)), it is a straightforward exercise to show that the expression for $\frac{S}{|Q|^{3/2}}$ given above is valid only to leading order in $J_{\phi}$ and $J_{\psi}$. The above result assumed $J_{\phi}J_{\psi}\geq 0.$ When $J_{\phi}J_{\psi}<0$, we get instead that
@@ 3.3 Fully nonlinear results
which was computed directly from Eqs. (14i), giving reassurance that our numerical code is correct.
@@ 4.1 Perturbative in the angular momenta
$Q_{3}(x,\rho)=\frac{\sigma_{1}(x)}{B_{2}}\left(1+\epsilon\,\rho^{\gamma}\,q_{3}(x)\right)$
$Q_{4}(x,\rho)=\frac{\sigma_{2}(x)}{B_{1}}\left(1+\epsilon\,\rho^{\gamma}\,q_{4}(x)\right)$
@@ 4.3 Fully nonlinear results
We now briefly discuss the $\lambda\neq 0$ case. For $\lambda=1$, all the scaling dimensions become integers. This is to be expected, since a known solution exists and is analytic in $\rho$ when changed to the relevant gauge (see Eq. (14i) for the corresponding near-horizon geometry). For $\lambda>1$ we always find negative scaling dimensions for at least one mode. These connect to the modes generated by ${\rm G}_{3}$ discussed in section 4.1. For $10<\lambda<$ 1 more interesting dynamics takes place. In particular, the orange and green families displayed in Fig. 2 both have negative or complex scaling dimensions, showing that these are not suitable to connect to an asymptotically flat end.
@@ 6 Discussion
Another surprise concerns the limit of vanishing angular momenta. We found that this limit depends on the ratio $J_{\phi}/J_{\psi}$ as both angular momenta are scaled to zero. One approaches the 5D Reissner-Nordström solution only when $J_{\phi}/J_{\psi}=1$. In general, one approaches a new family of static, extremal black holes which are not spherically symmetric. However these limiting solutions might have a stronger singularity on the horizon than the ones with $J_{i}\neq 0$. Linear perturbations of the near horizon geometry no longer scale to zero like a power law $\rho^{\gamma}$ (with $\gamma$ given in Eq. (45)) but can now diverge like $\log\rho$. If these perturbations are present in the full asymptotically flat solution, the metric would no longer be continuous on the horizon, and would not have a well defined horizon geometry. This deserves further investigation.